# Omnigenous umbilic stellarators

### v1

@@ Abstract
To better understand the dependence of the magnetic field structure in the plasma edge due to plasma boundary shape, we define and develop umbilic stellarators. These equilibria are characterized by a single continuous high-curvature edge on the plasma boundary that goes around multiple times toroidally before meeting itself. We develop a technique that allows us to simultaneously optimize the plasma boundary along with a curve lying on the boundary on which we impose a high curvature while imposing omnigenity — a property of the magnetic field that ensures trapped particle confinement throughout the plasma volume. After generating omnigenous umbilic stellarators, we design coil sets for some of these equilibria and calculate and understand the field line structure in the edge. Finally, we propose an experiment to modify an existing tokamak to a stellarator using this technique and explore a potential way to convert a limited tokamak into a diverted stellarator.
@@ 1 Introduction
Non-resonant divertors use a more complicated design to create a stochastic field in the edge, but similar to resonant divertors, they can be used to spread the load and improve plasma exhaust rejection. These divertors are also more resilient to changes in the magnetic field structure (Bader et al., 2017), but the magnetic field structure in the chaotic field region is complicated to calculate and requires further understanding. Recent work (Boozer & Punjabi, 2018; Punjabi & Boozer, 2022; Garcia et al., 2023; Davies et al., 2025) has further explored the possibility of using non-resonant divertors in stellarators.
Motivated by the need to understand how boundary shaping affects fieldline structure in the edge and divertor design, in this paper, we study the umbilic stellarator (US) as a method of achieving control over the edge magnetic field topology, and generate umbilic stellarators for vacuum and finite-$\beta$ cases111where $\beta=2\mu_{0}p/B^{2}$ is the ratio of the plasma to the magnetic pressure. Umbilic stellarators are characterized by regions of high curvature on the boundary and how these high-curvature regions are positioned. We develop a technique that allows us to obtain umbilic stellarators by simultaneously optimizing a curve and the plasma boundary, and imposing a high curvature along this curve which creates a ridge along the plasma boundary. Additionally, we ensure that all equilibria have reasonable omnigenity, by ensuring a low neoclassical transport. Finally, we propose an experiment to modify an existing tokamak to a stellarator using this technique and build it using umbilic coils. We show how an umbilic coil, different from helical coils, can create an umbilic edge in finite-pressure equilibria, acting as an alternate way to create a ridge on the boundary.
In §2, we explain the ideal MHD equilibrium. In §3, we introduce the umbilic boundary shape, how we approximate them with DESC, and how we solve the ideal MHD equations throughout the volume. In §4, we present a set of optimized vacuum($\beta=0$) and finite-$\beta$ stellarators and analyze their properties such as omnigenity. In §5, we take the Columbia HBT-EP plasma shape, modify it into an umbilic stellarator, and design electromagnetic coils for it. In §6, we conclude our work and discuss how it can be extended.
@@ 2 Ideal MHD equilibrium
where the plasma current $\boldsymbol{j}=(\boldsymbol{\nabla}\times\boldsymbol{B})/\mu_{0}$ from Ampere’s law, $p$ is the plasma pressure and $\mu_{0}$ is the vacuum magnetic permeability. Unlike an axisymmetric case, for stellarators we have to solve (3) as an optimization problem. We achieve this with the DESC (Dudt & Kolemen, 2020; Panici et al., 2023; Conlin et al., 2023; Dudt et al., 2023) stellarator optimization suite through minimization of $F\equiv\boldsymbol{j}\times\boldsymbol{B}-\boldsymbol{\nabla}p$. DESC can simultaneously solve an equilibrium while optimizing for multiple objectives such as MHD stability, quasisymmetry and more, without re-solving the equilibrium force balance equation (3) for each optimization variable. To reduce the computational cost further, we also use the fact that stellarators possess a discrete toroidal symmetry such that every stellarator comprises several identical sections. The number of identical sections is denoted by the field period $n_{\rm{FP}}$. Therefore, the force balance problem is only solved in a single field period.
In the following sections, we utilize DESC to optimize for omnigenous umbilic stellarators.
@@ 3 Umbilic stellarators
Umbilic stellarators draw inspiration from an umbilic bracelet (or torus) (Zeeman, 1976), a 3D shape with a cross-section that is a deltoid curve, a three-cusped hypocycloid which rotates by $\Delta\theta=2\pi/3$ poloidally after a single toroidal turn $\Delta\zeta=2\pi$. The sharp edge lies exactly on the cusp. Therefore, the sharp edge has to go around three times toroidally before meeting itself. A stellarator generated using an umbilic boundary is potentially advantageous for divertor designs, offering a continuous divertor option with a long connection length. For a smooth magnetic field, the sharp edge of such a boundary would be impossible for a magnetic field line. Therefore, the fieldline pitch on the boundary matches the umbilic topology That is, $\iota=n_{\rm{FP}}m/n=1/3$, where $m,n\in\mathbb{Z}$ corresponding to the poloidal and toroidal motion of the fieldline and $n_{\rm{FP}}$, the field period, is the measure of discrete symmetry of the stellarator. For a general umbilic shape, $m/n$ can attain rational values besides $1/3$. Further details are provided in this section.
@@ 3.1 Defining an umbilic surface and edge
where the major radius $R_{0}=1$ throughout this paper, unless stated otherwise, $\varrho$ is the minor radius of the umbilic-torus-like surface, $\theta$ is the poloidal angle used in DESC, $\zeta$ is the cylindrical toroidal angle, and $\lfloor$..$\rfloor$ … ⌋ is the floor function, defined such that
@@ 3.2 Approximating umbilic shapes with DESC
where $l,m$, and $n$ are the radial, poloidal, and toroidal mode numbers whereas $L,M$, and $N$ define the largest values of $l,m$, and $n$, respectively and $\mathcal{Z},\hat{F}$ are Zernike and Fourier basis functions. These parameters define the resolution of a DESC equilibrium.
@@
The sharpness of the umbilic edge increases as we increase the resolution of the fit surface and the agreement between DESC and the umbilic parametrization improves. Alternatively, one may use a double Chebyshev (Fortunato, 2024) instead of a double Fourier representation to improve the spectral fit, local bases methods such as finite-element to represent and solve the equilibrium, or use a non-linear transformation to map the circular torus in Fourier-Zernike basis to an umbilic shape, similar to the transformation used for circular tokamaks by Webster & Gimblett (2009). However, using a new basis and performing equilibrium optimization is beyond the scope of this work.
In the following section, we discuss how beginning with the parametrization of the umbilic surface, we can optimize for omnigenous umbilic stellarators.
@@ 3.3 Optimizing an umbilic equilibrium with DESC
where $\psi_{\mathrm{b}0}$ is the user-provided enclosed toroidal flux by the boundary and $\hat{\boldsymbol{p}}\in\{R_{\mathrm{b},mn},Z_{\mathrm{b},mn},a_{k}\}$ comprises parameters that determine the boundary shape and the position of the umbilic curve on the boundary. With the objective $\mathcal{F}_{\mathrm{stage-one}}$ defined, we run DESC on a single NVIDIA A100 GPU. A single optimization takes less than $90$ minutes.
@@ 4.1 US131($n_{\rm{FP}}=1,n=3,m=1$) omnigenous vacuum stellarator
To ensure that the high-curvature umbilic edge is aligned with the magnetic field lines, the boundary undergoes deformation, which results in an increased rotational transform. The optimized design minimized the objective $\mathcal{F}_{\mathrm{stage-one}}$ while ensuring that the normalized second principal curvature is the lowest ($\kappa_{2,\rho}<-35\rm{m}^{-1}$) along the umbilic edge. This causes the magnetic field line to stay close to the umbilic edge throughout the range $\zeta\in[0,6\pi]$ as shown in figure 3(c). In the limit where the edge is perfectly sharp, the field line will coincide with the umbilic edge.
The optimized US case is able to achieve an effective ripple $\epsilon_{\mathrm{eff}}$ of less than $2\%$ ($\epsilon_{\mathrm{eff}}^{3/2}<0.0028$) for most of the volume $(\rho<0.86)$. This ensures that neoclassical transport in the $1/\nu$ low-collisionality regime will not deteriorate the device performance. An interesting observation is that we can achieve a low ripple transport despite the lack of a specific helicity of the $B$ contours in the Boozer plot figure $4\mathit{(b)}$. This is an instance of a piecewise omnigenous (Velasco et al., 2024) field that is far from quasisymmetry. We prove this by analyzing the normalized second adiabatic invariant $\tilde{\mathcal{J}}_{\parallel}$ in Appendix C.
@@ 4.2 Modular coil design for the US131 equilibrium
on the plasma boundary. Note that in stage-two optimization, we ignore the surface current $\boldsymbol{K}=(\boldsymbol{B}_{\rm{out}}-\boldsymbol{B}_{\rm{in}})\times\boldsymbol{\hat{n}}$. Therefore, to verify our solution we calculate the field line trajectory due to the coils and create a Poincaré plot. The original fixed-boundary solution, its comparison with the Poincaré cross-section, and the optimized coilset are presented in figure 5.
@@
Modular coil design becomes complicated for umbilic equilibria as seen in the figure. The coils assume high curvature to reduce the normal field error along the umbilic edge. Increasing the number of coils improves the agreement of the fixed boundary solution with the Poincaré cross-section and reduces the max coil curvature but also reduces the coil-coil distance. However, the number of modular coils needed for these umbilic equilibria already exceeds the number of coils used for a single field period for a typical stellarator, which is almost always less than ten.
@@
Another important point is that even though the field lines do not exit in the regions of the highest curvature, the Poincaré traces stay aligned with the plasma boundary and flare near the umbilic edge, which still makes the umbilic edge a favorable location to put divertors or limiters.
[…]
@@
To ensure practical coil design for umbilic stellarators, a better technique would be to perform single-stage optimization, which simultaneously changes the plasma boundary and coil shapes, a union of $\mathcal{F}_{\mathrm{stage-one}}$ and $\mathcal{F}_{\mathrm{stage-two}}$, accounting for the coil complexity while maintaining omnigenity inside the plasma volume. Another possibility is to use umbilic coils, helical coils with a fractional helicity($m/n$) that create cusp-like regions, leading to a high curvature umbilic edge. We do a similar optimization to modify a tokamak in section 5, where we add an umbilic coil without changing the shape of the toroidal field coils to create an umbilic edge. In the next section, we shall use the same optimization process to create a finite-$\beta$, omnigenous umbilic stellarator.
@@ 4.3 US252 $(n_{\rm{FP}}=2,n=5,m=2)$ omnigenous finite-$\beta$ stellarator
In this section, we will generate a two field period $n=5,m=2$ finite-$\beta$ stellarator. We start with the umbilic parametrization defined in subsection 3.1 with $\varrho=0.1,r_{1}=0.133,r_{5}=0.1$ and use a pressure profile $p=5\times 10^{3}(1-\rho^{2})\mathrm{Pa}$, a rotational transform $\iota=0.755+0.043\rho^{2}$, and an enclosed toroidal flux $\psi_{\rm{b}}=\pi\hat{\rho}^{2}=0.00314\,\rm{T}m^{2}$. This creates an equilibrium with a $\beta=0.5\%$.
@@
Figure 6: Optimization output showing the initial and final magnetic field strengh Boozer plots in figures (a) and (b), respectively and the effective ripple in figure (c). The optimized configuration is another instance of piecewise omnigenity
We can see in figure 6, the ridges are only clearly visible in the initial equilibrium, which is close to quasisymmetry. The magnetic field of the optimized omnigenous equilibrium does not have any visible ridges even though the curvature along the umbilic ridge increases after optimization. This indicates that the appearance of ridges may only be a feature of quasisymmetric stellarators.
This equilibrium is optimized for poloidal omnigenity but, due to the umbilic curvature penalty, limits the poloidal omnigenity to the low-field region. The Boozer plot in figure 6(b) looks qualitatively similar to the piecewise-omnigenous configuration of the W7-X stellarator. The piecewise omnigenity of this equilibrium is demonstrated further in Appendix C. We also plot the cross-section of the equilibrium along with the second principal curvature in figure 7.
@@
We can see that the optimized equilibrium deforms and rotates as we move toroidally, while maintaining high curvature along the umbilic curve. The variation in the curvature is broad, indicating that imposing high curvature at specific locations may be more suitable than others. This can be readily accomplished by selecting sections of the umbilic curve where high curvature is applied and could be a potential way to place potential locations for non-resonant divertors (Bader et al., 2018b).

### v2

@@ Abstract
To better understand the dependence of the magnetic field structure in the plasma edge on the plasma boundary shape, in the context of X-point and island divertor designs, we define and develop a class of stellarators called umbilic stellarators. These equilibria are characterized by a single continuous high-curvature edge on the plasma boundary that goes around multiple times toroidally before meeting itself. We develop a technique that allows us to simultaneously optimize the plasma boundary along with a curve lying on the boundary on which we impose a high curvature while imposing omnigenity — a property of the magnetic field that ensures trapped particle confinement throughout the plasma volume. We find that umbilic stellarators naturally tend to favor piecewise omnigenity instead of omnigenity with a specific helicity. After generating omnigenous umbilic stellarators, we design coil sets for some of them and explore the field line structure in the edge and its sensitivity to small fluctuations in the plasma. Finally, using single-stage optimization, we simultaneously modify the plasma and coil shape and propose an experiment to modify an existing tokamak to a finite-$\beta$ stellarator using this technique and explore a potentially simpler way to convert a limited tokamak into a diverted stellarator.
@@ 1 Introduction
Non-resonant divertors use a more complicated design to create a stochastic field in the edge, but similar to resonant divertors, they can be used to spread the heat load and improve plasma exhaust rejection. These divertors are also more resilient to changes in the magnetic field structure (Bader et al., 2017), but the magnetic field structure in the chaotic field region is complicated to calculate and requires further understanding. Recent work (Boozer & Punjabi, 2018; Punjabi & Boozer, 2022; Garcia et al., 2023; Davies et al., 2025) has further explored the possibility of using non-resonant divertors in stellarators.
Motivated by the need to understand how boundary shaping affects fieldline structure in the edge and X-point and island divertor design, in this paper, we study the umbilic stellarator (US) as a method of achieving control over the edge magnetic field topology, and generate umbilic stellarators for vacuum and finite-$\beta$ cases111where $\beta=2\mu_{0}p/B^{2}$ is the ratio of the plasma to the magnetic pressure. Umbilic stellarators are characterized by regions of high curvature on the boundary and how these high-curvature regions are positioned. We develop a technique that allows us to obtain umbilic stellarators by simultaneously optimizing a curve and the plasma boundary, and imposing a high curvature along this curve which creates a ridge along the plasma boundary. Additionally, we ensure that all equilibria have reasonable omnigenity, ensuring low neoclassical transport. Finally, we propose an experiment to modify an existing tokamak to a stellarator using our optimization process and build it using umbilic coils, which are helically wound coils, similar to divertor coils, with fractional helicity that carry a small current (compared to the plasma current), and follow an umbilic edge on the plasma boundary. We show how an umbilic coil can create a high-curvature umbilic region in finite-pressure equilibria, acting as an alternative to modular coils, to create a ridge on the plasma boundary.
In §2, we explain the ideal MHD equilibrium. In §3, we introduce the umbilic boundary shape, how to parametrize them, and how we approximate them with DESC, and how to solve the ideal MHD equations throughout the volume enclosed by the approximated shape. In §4, we present a set of optimized vacuum($\beta=0$) and finite-$\beta$ stellarators and analyze their properties such as omnigenity. To understand the magnetic field structure in the edge, we design modular coils for the vacuum equilibrium and trace the fieldlines throughout the volume. In order to assess the robustness of the edge magnetic field configuration, we introduce a dummy current source at the magnetic axis to emulate a plasma fluctuation. In §5, we take the Columbia HBT-EP plasma shape, modify it into an umbilic stellarator without changing the coilset used for the original tokamak, just by adding an umbilic coil. We study cases where the umbilic coil carries current in the same (co-$I$) and opposite (counter-$I$) direction as the plasma current $I$, study the effect of the umbilic current on edge iota, and solve a single stage problem for the co-$I$ case, by simultaneously modifying the plasma boundary and umbilic coil shape. In §6, we conclude our study by highlighting existing limitations and exploring potential avenues for overcoming them.
@@ 2 Ideal MHD equilibrium
where the plasma current $\boldsymbol{j}=(\boldsymbol{\nabla}\times\boldsymbol{B})/\mu_{0}$ from Ampere’s law, $p$ is the plasma pressure and $\mu_{0}$ is the vacuum magnetic permeability. Unlike an axisymmetric case, for stellarators we have to solve (3) as an optimization problem. We achieve this with the DESC (Dudt & Kolemen, 2020; Panici et al., 2023; Conlin et al., 2023; Dudt et al., 2023) stellarator optimization suite through minimization of $F\equiv\boldsymbol{j}\times\boldsymbol{B}-\boldsymbol{\nabla}p$. DESC can simultaneously solve an equilibrium while optimizing for multiple objectives such as MHD stability, quasisymmetry and more, without re-solving the equilibrium force balance equation (3) for each optimization variable. To reduce the computational cost further, we also use the fact that stellarators possess a discrete toroidal symmetry such that every stellarator comprises several identical sections. The number of identical sections is denoted by the field period $n_{\rm{FP}}$. Therefore, the force balance problem is only solved in a single field period. In the following sections, we utilize DESC to optimize for omnigenous umbilic stellarators.
@@ 3 Umbilic stellarators
Umbilic stellarators draw inspiration from an umbilic bracelet (or torus) (Zeeman, 1976), a 3D shape with a cross-section that is a deltoid curve, a three-cusped hypocycloid which rotates by $\Delta\theta=2\pi/3$ poloidally after a single toroidal turn $\Delta\zeta=2\pi$. The sharp edge lies exactly on the cusp. Therefore, the sharp edge has to go around three times toroidally before meeting itself. A stellarator generated using an umbilic boundary is potentially advantageous for divertor designs, offering a continuous divertor option with a long connection length. For a smooth magnetic field, the sharp edge of such a boundary would be impossible for a magnetic field line. Therefore, the fieldline pitch on the boundary matches the umbilic topology. That is, $\iota=n_{\rm{FP}}m/n=1/3$, where $m,n\in\mathbb{Z}$ corresponding to the poloidal and toroidal motion of the fieldline and $n_{\rm{FP}}$, the field period, is the measure of discrete symmetry of the stellarator. For a general umbilic shape, $m/n$ can attain rational values besides $1/3$. Further details are provided in this section.
@@ 3.1 Defining an umbilic surface and edge
where the major radius $R_{0}=1$ throughout this paper, unless stated otherwise, $\varrho$ is the minor radius of the umbilic-torus-like surface, $\theta$ is the poloidal angle used in DESC, $\zeta$ is the cylindrical toroidal angle, and $\lfloor$$\rfloor$ is the floor function, defined such that
@@ 3.2 Approximating umbilic shapes with DESC
where $l,m$, and $n$ are the radial, poloidal, and toroidal mode numbers, whereas $L,M$, and $N$ define the largest values of $l,m$, and $n$, respectively, and $\mathcal{Z},\hat{F}$ are Zernike and Fourier basis functions. These parameters define the resolution of aDESC equilibrium.
@@
The sharpness of the umbilic edge increases as we increase the resolution of the fit surface and the agreement between DESC and the umbilic parametrization improves. Alternatively, one may use a double Chebyshev (Fortunato, 2024) instead of a double Fourier representation to improve the spectral fit, local basis methods such as finite-element to represent and solve the equilibrium, or use a non-linear transformation to map the circular torus in Fourier-Zernike basis to an umbilic shape, similar to the transformation used for circular tokamaks by Webster & Gimblett (2009). However, using a new basis and performing equilibrium optimization is beyond the scope of this work.
In the following section, we discuss how, beginning with the parametrization of the umbilic surface, we can optimize for omnigenous umbilic stellarators.
@@ 3.3 Optimizing an umbilic equilibrium with DESC
[…]
@@ 4.1 US131($n_{\rm{FP}}=1,n=3,m=1$) omnigenous vacuum stellarator
To ensure that the high-curvature umbilic edge is aligned with the magnetic field lines, the boundary undergoes deformation, which results in an increased rotational transform. The optimized design minimized the objective $\mathcal{F}_{\mathrm{stage-one}}$ while ensuring that the normalized second principal curvature is the lowest along the umbilic edge. The curvature on the umbilic edge ranges from $\kappa_{2,\rho}\in[-187,-70]\,m^{-1}$ whereas the average curvature on the rest of the boundary is $\mathrm{avg}(\kappa_{2,\rho})=-12\,m^{-1}$. This causes the magnetic field line to stay close to the umbilic edge throughout the range $\zeta\in[0,6\pi]$ as shown in figure 3(c). In the limit where the edge is perfectly sharp, the field line will coincide with the umbilic edge.
The optimized US case is able to achieve an effective ripple $\epsilon_{\mathrm{eff}}$ of less than $2\%$ ($\epsilon_{\mathrm{eff}}^{3/2}<0.0028$) for most of the volume $(\rho<0.86)$ and $\epsilon_{\mathrm{eff}}^{3/2}<=0.001$ for $\rho<=0.6$. This ensures that neoclassical transport in the $1/\nu$ low-collisionality regime will not deteriorate the device performance in the core but may deteriorate confinement in the edge. An interesting observation is that we can achieve a low ripple transport despite the lack of a specific helicity of the $B$ contours in the Boozer plot figure $4\mathit{(b)}$. This is an instance of a piecewise omnigenous (Velasco et al., 2024) field that is far from quasisymmetry. Since we did not seek omnigenity with a specific helicity ($w_{\mathrm{O}}=0$), the effective ripple penalty acted as a proxy for omnigenity, leading to a piecewise-omnigenous configuration. To prove that these configurations are piecewise omnigenous, we analyze the normalized second adiabatic invariant $\tilde{\mathcal{J}}_{\parallel}$ in Appendix C.
@@ 4.2 Modular coil design for the US131 equilibrium
on the plasma boundary. Note that in stage-two optimization, we ignore the surface current $\boldsymbol{K}=(\boldsymbol{B}_{\rm{out}}-\boldsymbol{B}_{\rm{in}})\times\boldsymbol{\hat{n}}$. Therefore, to verify our solution, we calculate the field line trajectory due to the coils and create a Poincaré plot. The original fixed-boundary solution, its comparison with the Poincaré cross-section, and the optimized coilset are presented in figure 5.
@@
Note that even though the curvature along the umbilic ridge is high, it does not behave like a perfect X-point (or X-line) as the fieldline structure becomes chaotic in the edge and resembles the “turnstiles” topology discussed by Punjabi & Boozer (2022). However, even though the field lines do not exit in the regions of the highest curvature, like a tokamak X-point, the Poincaré traces of the fieldlines stay aligned with the plasma boundary and flare near the umbilic edge, which still makes the umbilic edge a favorable location to place divertors or limiters. This study indicates that the behavior of magnetic field lines for a high-curvature ridge can differ significantly from a perfectly sharp edge.
Another observation is that modular coil design becomes complicated for umbilic equilibria as seen in the figure. The coils assume high curvature to reduce the normal field error along the umbilic edge. Increasing the number of coils improves the agreement of the fixed boundary solution with the Poincaré cross-section and reduces the max coil curvature but also reduces the coil-coil distance. However, the number of modular coils needed for these umbilic equilibria already exceeds the number of coils used for a single field period for a typical stellarator, which is almost always less than ten.
Finally, we note that the regions of maximum normal field error are parallel to the umbilic edge so that, on average, $\boldsymbol{b}\cdot\boldsymbol{\nabla}(\boldsymbol{B}_{\mathrm{out}}\cdot\hat{\boldsymbol{n}})=0$ [^3] We use the magnetic field unit vector instead of the vector along the umbilic edge because the magnetic field is reasonably aligned with the umbilic edge throughout its span as shown in figure 3(c). This pattern indicates a hidden relationship between the normal field error and the second principal curvature in umbilic stellarators.
@@
To ensure practical coil design for umbilic stellarators, a better technique would be to perform single-stage optimization, which simultaneously changes the plasma boundary and coil shapes, a union of $\mathcal{F}_{\mathrm{stage-one}}$ and $\mathcal{F}_{\mathrm{stage-two}}$, accounting for the coil complexity while maintaining omnigenity inside the plasma volume. This is done in §5.2 for a simpler problem. Another possibility is to use umbilic coils, helically wound divertor-like coils with a fractional helicity($m/n$) that carry a small current that pushes (pulls) the plasma away from (towards) itself, creating cusp-like regions, leading to a high curvature umbilic edge. We do a similar optimization to modify a tokamak in section §5.1, where we add an umbilic coil without changing the shape of the toroidal field coils to create an umbilic edge. In the next section, we shall use the same optimization process to create a finite-$\beta$, omnigenous umbilic stellarator.
4.3 US131 sensitivity analysis of the field structure in the edge
In this section, we shall introduce a line current source along the magnetic axis of the US131 configuration. The line current aims to simulate an internal fluctuation within the plasma. For different values of the current source strength, we will perform the field line integration across the entire volume. The purpose of this investigation is to assess the resilience of the field line structure at the edge of the US131 stellarator and to determine the extent of modification required for the divertor configuration. We present the fieldline structure for different values of the current in figure 6.
(a) $I_{\mathrm{axis}}=250A$
(b) $I_{\mathrm{axis}}=500A$
(c) $I_{\mathrm{axis}}=1000A$
Figure 6: Poincare plots at $\zeta=0$ with a current source placed on-axis for different values of the current, compared with the original fixed boundary solution (in red). Increasing the current significantly reduces the nestedness of the flux surfaces but the fieldline structure in the edge still “flares” out near the high-curvature edges and the divertor placement does not need to be modified.
We observe that even though the flux surfaces inside the plasma boundary lose nestedness and become stochastic, the divertor structure used for the unperturbed US131 case is still an effective choice for this equilibrium, i.e., the fieldline structure in the edge is resilient to fluctuations inside the plasma. Another observation is the emergence of “O”-point-like structures near the high-curvature regions in figure 6(b).
@@ 4.4 US252 $(n_{\rm{FP}}=2,n=5,m=2)$ omnigenous finite-$\beta$ stellarator
In this section, we will generate a two-field period $n=5,m=2$ finite-$\beta$ stellarator. We start with the umbilic parametrization defined in subsection 3.1 with $\varrho=0.1,r_{1}=0.133,r_{5}=0.1$ and use a pressure profile $p=5\times 10^{3}(1-\rho^{2})\mathrm{Pa}$, a rotational transform $\iota=0.755+0.043\rho^{2}$, and an enclosed toroidal flux $\psi_{\rm{b}}=\pi\hat{\rho}^{2}=0.00314\,\rm{T}m^{2}$. This creates an equilibrium with a $\beta=0.5\%$.
@@
Figure 7: Optimization output showing the initial and final magnetic field strength Boozer plots in figures (a) and (b), respectively and the effective ripple in figure (c). The black lines (a) correpond to the magnetic field lines and the dashed lines correspond to the umbilic curve. in The optimized configuration is another instance of piecewise omnigenity
We can see in figure 7, the ridges in the Boozer plots are only clearly visible in the initial equilibrium, which is close to quasisymmetry. The magnetic field of the optimized omnigenous equilibrium does not have any visible ridges in the Boozer plot, even though the curvature along the umbilic ridge increases after optimization. This indicates that the appearance of ridges in the Boozer representation may only be a feature of quasisymmetric stellarators.
This equilibrium is optimized for poloidal omnigenity but, due to the umbilic curvature penalty, limits the poloidal omnigenity to the low-field region. The Boozer plot in figure 7(b) looks qualitatively like a piecewise-omnigenous Velasco et al. (2024) configuration, which looks similar to the magnetic field in the W7-X stellarator. The piecewise omnigenity of this equilibrium is demonstrated further in Appendix C. We also plot the cross-section of the equilibrium along with the second principal curvature in figure 8. The neoclassical transport coefficient $\epsilon_{\mathrm{eff}}^{3/2}<0.001$ for most ($\rho<0.7$) of the volume, which means the equilibrium is omnigenous in the core.
@@
We also observe that the optimized equilibrium deforms and rotates as we move toroidally, while maintaining high curvature along the umbilic curve. The curvature along the umbilic edge $\kappa_{2,\rho}\in[-67,-36]\,m^{-1}$ while the average curvature on the rest of the plasma boundary is $\mathrm{avg}(\kappa_{2,\rho})=-8\,m^{-1}$. The variation in the curvature is broad, indicating that imposing high curvature at specific locations may be more suitable than others. This can be readily accomplished by selecting sections of the umbilic curve where high curvature is applied and could be a potential way to place potential locations for non-resonant divertors (Bader et al., 2018b).